# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v3

@@ A.2.2 Discussions on Radius Restrictions
By setting $\delta<\ \frac{a_{0}}{2\delta}$, we have $\sup_{\mathsf{P}_{i}\in\mathcal{P}}\tau((\mathsf{P}^{\pi}_{m}\cdots\mathsf{P}^{\pi}_{1}))<1$, and by Lemma A.3,
$\hat{\rho}\bigl(Q^{\pi}_{\mathcal{P}}\bigr)\ \leq\ \Bigl(sup_{\mathsf{P}_{i}\in\mathcal{P}}\tau((\mathsf{P}^{\pi}_{m}\cdots\mathsf{P}^{\pi}_{1}))\Bigr)^{1/m}\ \leq\ \bigl(1-(a_{0}-2m\delta)\bigr)^{1/m}\ <\ 1.$ (68)
@@ A.2.3 Extremal Norm Construction
$\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}\leq\alpha^{-k}C\lambda^{k}\|x\|_{2}=C\left(\frac{\lambda}{\alpha}\right)^{-k}\|x\|_{0}\longrightarrow 0\quad\text{as}\quad k\rightarrow\infty$ (76)
@@
$\|x+y\|_{\rm est}=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(x+y)\|_{2}\leq\|x\|_{\rm est}+\|y\|_{\rm est}$ (79)
@@ Appendix B Biased Stochastic Approximation Convergence Rate
By [42, Proposition 1–2], using the Moreau envelope function $M(x)$ in Definition 2.2 of [7], we define
@@ Appendix C Uncertainty Set Support Function Estimators
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{36(1+\frac{1}{\delta})^{2}2^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (146)
@@
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2\sum_{n=0}^{N_{\max}}\frac{362^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (147)
@@ Appendix D Convergence for Robust TD
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (154)
@@
$\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}$ (155)
@@
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6\cdot 2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (156)
@@
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$ $\leq\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304t_{\mathrm{mix}}^{2}N_{\max}$ (157)
@@
$\varepsilon^{\text{TV}}_{\text{bias}}=48C_{\mathcal{P}}(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (160)
@@
$A^{\text{TV}}$ $=2C_{\mathcal{P}}^{2}\left(24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}(1+\frac{1}{\delta})^{2}t_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (161)
@@
$\varepsilon^{\text{Wass}}_{\text{bias}}=48C_{\mathcal{P}}\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (162)
@@
$A^{\text{Wass}}$ $=2C_{\mathcal{P}}^{2}\left(24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}N_{\max}\right)$ (163)
@@
We use the result from Theorem D.2, to set $\mathbb{E}\Bigl[\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr]\leq\epsilon^{2}$. For contamination uncertainty set we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 5.1, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
@@
$+\frac{16\sqrt{2A^{\mathrm{Cont}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\log T}{\sqrt{T}(1-\gamma)}\right)$ (183)
@@
$+\frac{16\sqrt{2A^{\mathrm{TV}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{N_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right)$ (184)
@@
$+\frac{16\sqrt{2A^{\mathrm{Wass}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{N_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right)$ (185)
Where all variables are defined the same as in Theorem D.2.
@@
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (190)
@@
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (191)
@@
We use the result from Theorem D.3, to set $\mathbb{E}\Bigl[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr]\leq\epsilon$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{3}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 5.1, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{4}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.

### v4

@@ A.2.2 Discussions on Radius Restrictions
By setting $\delta<\ \frac{a_{0}}{2m}$, we have $\sup_{\mathsf{P}_{i}\in\mathcal{P}}\tau((\mathsf{P}^{\pi}_{m}\cdots\mathsf{P}^{\pi}_{1}))<1$, and by Lemma A.3,
$\hat{\rho}\bigl(Q^{\pi}_{\mathcal{P}}\bigr)\ \leq\ \Bigl(\sup_{\mathsf{P}_{i}\in\mathcal{P}}\tau((\mathsf{P}^{\pi}_{m}\cdots\mathsf{P}^{\pi}_{1}))\Bigr)^{1/m}\ \leq\ \bigl(1-(a_{0}-2m\delta)\bigr)^{1/m}\ <\ 1.$ (68)
@@ A.2.3 Extremal Norm Construction
$\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}x\|_{2}\leq\alpha^{-k}C\lambda^{k}\|x\|_{2}=C\left(\frac{\lambda}{\alpha}\right)^{-k}\|x\|_{2}\longrightarrow 0\quad\text{as}\quad k\rightarrow\infty$ (76)
@@
$\|x+y\|_{\rm ext}=\sup_{k\geq 0}\sup_{Q_{1},\ldots,Q_{k}\in Q^{\pi}_{\mathcal{P}}}\alpha^{-k}\|Q_{k}Q_{k-1}\ldots Q_{1}(x+y)\|_{2}\leq\|x\|_{\rm ext}+\|y\|_{\rm ext}$ (79)
@@ Appendix B Biased Stochastic Approximation Convergence Rate
By [45, Proposition 1–2], using the Moreau envelope function $M(x)$ in Definition 2.2 of [8], we define
@@ Appendix C Uncertainty Set Support Function Estimators
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2S\sum_{n=0}^{N_{\max}}\frac{36(1+\frac{1}{\delta})^{2}2^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}S\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (146)
@@
$\mathbb{E}\left[\left(\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right)^{2}\right]$ $\leq 2\|V\|_{\mathrm{sp}}^{2}+2S\sum_{n=0}^{N_{\max}}\frac{362^{-n}\|V\|_{\mathrm{sp}}^{2}}{2^{-(n+1)}}$
$=2\|V\|_{\mathrm{sp}}^{2}+144S\|V\|_{\mathrm{sp}}^{2}N_{\max}$ (147)
@@ Appendix D Convergence for Robust TD
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (154)
@@
$\leq\left(24(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144(1+\frac{1}{\delta})^{2}S\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304(1+\frac{1}{\delta})^{2}St_{\mathrm{mix}}^{2}N_{\max}$ (155)
@@
$\left\lVert\mathbb{E}\left[\hat{\mathbf{T}}_{g_{0}}(V_{t})\right]-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert_{\infty}\leq 6\sqrt{S}2^{-\frac{N_{\max}}{2}}\|V\|_{\mathrm{sp}}=24\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (156)
@@
$\mathbb{E}\left[\left\lVert\hat{\mathbf{T}}_{g_{0}}(V_{t})-{\mathbf{T}}_{g_{0}}(V_{t})\right\rVert^{2}_{\infty}\right]$ $\leq\left(24\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+3\|V\|_{\mathrm{sp}}^{2}+144\|V\|_{\mathrm{sp}}^{2}N_{\max}$
$\leq\left(24\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+48t_{\mathrm{mix}}^{2}+2304St_{\mathrm{mix}}^{2}N_{\max}$ (157)
@@
$\varepsilon^{\text{TV}}_{\text{bias}}=48C_{\mathcal{P}}(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (160)
@@
$A^{\text{TV}}$ $=2C_{\mathcal{P}}^{2}\left(24(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}(1+\frac{1}{\delta})^{2}St_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}SN_{\max}\right)$ (161)
@@
$\varepsilon^{\text{Wass}}_{\text{bias}}=48C_{\mathcal{P}}\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}=\operatorname*{\mathcal{O}}\left(\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)$ (162)
@@
$A^{\text{Wass}}$ $=2C_{\mathcal{P}}^{2}\left(24\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\right)^{2}+96C_{\mathcal{P}}^{2}t_{\mathrm{mix}}^{2}+4608C_{\mathcal{P}}^{2}St_{\mathrm{mix}}^{2}N_{\max}$
$=\operatorname*{\mathcal{O}}\left(t_{\mathrm{mix}}^{2}SN_{\max}\right)$ (163)
@@
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{St^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{St^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (172)
@@
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{St^{2}_{\mathrm{mix}}N_{\max}}{T(1-\gamma)^{2}}+\frac{St^{2}_{\mathrm{mix}}2^{-\frac{N_{\max}}{2}}\log T}{(1-\gamma)^{2}}\right)$ (174)
@@
We use the result from Theorem D.2, to set $\mathbb{E}\Bigl[\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr]\leq\epsilon^{2}$. For contamination uncertainty set we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{\sqrt{S}t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{\sqrt{S}t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 5.1, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{\sqrt{S}t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
@@
$+\frac{16\sqrt{2A^{\mathrm{Cont}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\log T}{\sqrt{T}(1-\gamma)}\right).$ (183)
@@
$+\frac{16\sqrt{2A^{\mathrm{TV}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48(1+\frac{1}{\delta})\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{SN_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{\sqrt{S}t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right).$ (184)
@@
$+\frac{16\sqrt{2A^{\mathrm{Wass}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\sqrt{SN_{\max}}\log T}{\sqrt{T}(1-\gamma)}+\frac{\sqrt{S}t_{\mathrm{mix}}2^{-\frac{N_{\max}}{4}}\log^{\frac{3}{2}}T}{\sqrt{1-\gamma}}+\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T\right).$ (185)
where all the above variables are defined the same as in Theorem D.2.
@@
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24\sqrt{S}(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (190)
@@
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|\leq 24\sqrt{S}2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}$ (191)
@@
We use the result from Theorem D.3, to set $\mathbb{E}\Bigl[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr]\leq\epsilon$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{\sqrt{S}t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{3}\frac{\sqrt{S}t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 5.1, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{4}\frac{\sqrt{S}t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.